# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v1

@@ 5 Analysis
To isolate the impact of initialization on training, we perform truncated SVD on various matrices, including Kaiming initialization and $\Delta W_{avg}$ with varying levels of added Gaussian noise, as shown in Table 4. By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization. The results clearly demonstrate the significance of initialization—our approach consistently outperforms other variants.
@@
We aim to examine the effect of optimal gradient approximation on training. Specifically, we want $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ without enforcing $B^{T}B=AA^{T}=I$. We achieve this through the following steps:
@@
This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$. This setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3.
We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3. Although both methods start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation. Furthermore, LoRA-SB achieves higher accuracies on the GSM8K and MATH benchmarks, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
@@
Training Time Overhead vs LoRA-XS
As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset. Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS. The results show that the additional overhead is negligible, adding just 242–4 minutes compared to the total training time of 353–5 hours per epoch ($\approx 1.1\%$ to $1.3\%$). Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
@@ 6 Conclusion
In this work, we introduced LoRA-SB, a novel framework that bridges the gap between low-rank parameter-efficient fine-tuning and full fine-tuning. This is enabled by our carefully designed initialization strategy, which approximates the first step of full fine-tuning and ensures that the most relevant subspaces for task-specific adaptation are captured from the outset. Our theoretical analysis uncovered critical limitations in LoRA-XS: suboptimal gradient approximation and reliance on fixed initialization that fails to adapt to task-specific nuances. By addressing both issues through a unified approach, we demonstrate that LoRA-SB achieves optimal scaling and preserves critical update directions. Furthermore, our approach ensures hyperparameter independence by directly approximating the full fine-tuning gradient, thereby eliminating instability and convergence issues associated with scaling factors.
Through extensive experiments on 4 models across 16 datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method exceeds the performance of LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS. Our work advances the state of PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks. Future work includes exploring adaptive layer-wise rank settings and evaluating the approach on other architectures, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
@@ A.1 Proof of Lemma 1
Let $\Delta W$ be an update learned by LoRA-XS . Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\mathcal{W}_{LoRA-XS}:=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
@@
The column space of $\Delta W$ is given by:
$\text{Col}(\Delta W)$ $=\{y\in\mathbb{R}^{m}\mid y=\Delta Wx,\,x\in\mathbb{R}^{n}\}$ (20)
Now since $\Delta W=BRA$:
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\dots\text{ for }z=RAx$ (21)
Now:
$\text{Col}(B)$ $=\{y\in\mathbb{R}^{m}\mid y=Bx,\,x\in\mathbb{R}^{r}\}$ (22)
Now since $\text{Col}(RA)\subseteq\mathbb{R}^{r}$:
$\text{Col}(\Delta W)$ $\subseteq\text{Col}(B)$ (23)
A symmetric proof can now be given for row space of $\Delta W$. ∎
@@ A.2 Proof of Lemma 2
During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
@@
$g:=\frac{\partial L}{\partial W}\quad\&\quad g^{R}_{\text{LoRA-XS}}:=\frac{\partial L}{\partial R}$ (24)
Now by chain rule:
$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial R}\implies\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}\frac{\partial X}{\partial R}\dots\text{ for }X=RA$ (25)
@@
$\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}=sB^{T}g\implies\frac{\partial L}{\partial R}=sB^{T}g\frac{\partial X}{\partial R}$ (26)
Let $sB^{T}g=y$. We know that when $X=RA$:
$y\frac{\partial X}{\partial R}=yA^{T}\implies\frac{\partial L}{\partial R}=yA^{T}=sB^{T}gA^{T}\$(27)
$\text{Therefore, \quad}\boxed{g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}}$ (28)
@@ A.3 Proof of Theorem 3
[…]
@@
Since we already defined the equivalent gradient $\tilde{g}:=sBg^{R}A$, the minimization problem can be formally denoted as:
@@
$B^{T}gA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}sBg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}Bg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s^{2}$ (32)
Now since $B$ and $A$ are full rank, multiplying both sides by $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$ on the left and right side respectively gives:
$(B^{T}B)^{-1}(B^{T}Bg^{R}AA^{T})(AA^{T})^{-1}$ $=(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}/s^{2}$ (33)
$\text{Therefore,}\quad\boxed{g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}}$ (34)
@@ A.4 Proof of Theorem 4
Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate. This update guarantees a reduction in the loss function, similar to traditional gradient descent methods. Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta{\langle g^{R}_{LoRA-XS},g^{R}\rangle}=-\eta{\langle g^{R}_{LoRA-XS},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}\rangle}\leq 0$
@@
We establish that during optimization, the differential change in the loss function, $\mathrm{d}L$, can be expressed as:
$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ (35)
Let us derive this expression step by step. First, the differential change in loss can be written as:
$\mathrm{d}L=\langle\frac{\partial L}{\partial R},\mathrm{d}R\rangle_{F}$ (36)
The update equation for $R$ follows:
$\mathrm{d}R=-\eta g^{R}$ (37)
Given that $\frac{\partial L}{\partial R}=g^{R}_{\text{LoRA-XS}}$, we obtain:
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}\rangle_{F}$ (38)
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$
To prove that $\mathrm{d}L\leq 0$, we demonstrate that:
$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\geq 0$ (39)
We first establish that $(B^{T}B)^{-1}$ is positive definite. For any non-zero vector $x$, since $B$ is full-rank:
$\langle x,B^{T}Bx\rangle=\langle Bx,Bx\rangle=\|Bx\|^{2}>0$ (40)
This proves $B^{T}B$, and consequently, $(B^{T}B)^{-1}$ are positive definite. Through Cholesky decomposition, we can express $(B^{T}B)^{-1}=VV^{T}$. Similarly, $(AA^{T})^{-1}$ is positive-definite and can be decomposed as $(AA^{T})^{-1}=UU^{T}$.
Therefore:
$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{R}_{\text{LoRA-XS}},VV^{T}g^{R}_{\text{LoRA-XS}}UU^{T}\rangle_{F}$ (41)
$=\frac{1}{s^{2}}\langle V^{T}g^{R}_{\text{LoRA-XS}}U,V^{T}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
$=\frac{1}{s^{2}}\|V^{T}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0$
Thus, we have proven:
$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\leq 0$ (42)
For our specific initialization where $(B^{T}B)=I$, $(AA^{T})=I$, and $s=1$, the result simplifies to:
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}\leq 0$ (43)
@@ A.5 Proof of Theorem 5
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\dots\text{ but not when }\tilde{g}=sBg^{R}_{LoRA-XS}A$
@@
Let $g$ be the full fine-tuning gradient. We want to prove that $\tilde{g}$ does not depend on $s$, so we try to express it in terms of $g$ which does not depend on the LoRA-XS training process or reparameterization
@@
$g^{R}=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\implies\tilde{g}=\frac{s}{s^{2}}B(B^{T}B^{-1})g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}A$ (44)
Now since $g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}$:
$\tilde{g}=\frac{1}{s}B(B^{T}B^{-1})sB^{T}gA^{T}(AA^{T})^{-1}A=B(B^{T}B^{-1})B^{T}gA^{T}(AA^{T})^{-1}A$ (45)
which is $s$ independent!
@@
$g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}\implies\tilde{g}=sB(sB^{T}gA^{T})A\implies\tilde{g}=s^{2}BB^{T}gA^{T}A$ (46)
which is not $s$ independent! ∎
@@ A.6 Proof of Theorem 6
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
$\Delta R=-\eta\nabla_{R}\mathcal{L}(R)\implies B\Delta RA=-\eta B\nabla_{R}\mathcal{L}(R)A$ (47)
@@
$\Delta W$ $=-\eta\nabla_{W}\mathcal{L}(W_{0})$ (48)
@@
$\|B\Delta RA-\eta\mathcal{L}_{W}(W_{0})\|_{F}$ $=\eta\|B\nabla_{R}\mathcal{L}(R)A-\mathcal{L}_{W}(W_{0})\|_{F}$ (49)
@@
$\nabla_{R}\mathcal{L}$ $=B^{T}\nabla_{W}\mathcal{L}A^{T}$ (50)
The problem becomes:
$\min_{A_{\text{init}},B_{\text{init}}}\|B^{T}(B^{T}\nabla_{W}\mathcal{L}A^{T})A-\nabla_{W}\mathcal{L}\|_{F}\dots\text{ where }\nabla_{W}\mathcal{L}=USV^{T}$ (51)
@@
$\|BB^{T}\nabla_{W}\mathcal{L}A^{T}A-\nabla_{W}\mathcal{L}\|_{F}$ $=\|U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}-UsV^{T}\|_{F}$ (52)
Now:
$U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ $=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (53)
Now rank of $W^{\prime}$ where:
$W^{\prime}$ $=U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ (54)
is $\leq r$ since rank of $B_{\text{init}}$ and $A_{\text{init}}$ is $r$
By Eckart-Young Theorem, the optimal low-rank optimization w.r.t Frobenius norm is:
$W^{\prime*}$ $=\underset{\text{rank}(W^{\prime})=r}{\text{arg min}}\|W^{\prime}-\nabla_{W}\mathcal{L}\|_{F}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (55)
Since we get an identical expression, our solution is optimal. ∎
Appendix B Inference Costs

### v3

@@ 5 Analysis
To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (He et al., 2015) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4. By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization. The results clearly demonstrate the significance of initialization—our approach consistently outperforms other variants.
@@
We aim to examine the effect of optimal gradient approximation. Specifically, we want $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ without enforcing $B^{T}B=AA^{T}=I$. We achieve this through:
@@
This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$. The setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3. We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3. Although both start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation. Furthermore, LoRA-SB achieves higher accuracies on GSM8K and MATH, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
@@
Training Time Overhead vs LoRA-XS.
As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset. Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS. The results show that the additional overhead is negligible, adding just 24242–42 – 4 minutes compared to the total training time of 35353–53 – 5 hours per epoch ($\approx 1.1\%$ to $1.3\%$). Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
@@ 6 Conclusion
In this work, we introduced LoRA-SB, which bridges the gap between low-rank PEFT and full FT. This is enabled by our initialization strategy, which approximates the first step of full FT and ensures that the most relevant subspaces for task-specific adaptation are captured. We achieve optimal gradient scaling and preserve update directions throughout training. Our approach ensures hyperparameter independence by approximating the full FT gradient, thereby eliminating instability issues associated with scaling factors.
Through extensive experiments, we demonstrate that our method outperforms LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS. Our work advances PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks. Future work includes exploring adaptive layer-wise rank settings and integrating LoRA-SB with quantization. We also aim to evaluate its performance on other models, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
@@ A.1 Proof of Lemma 1
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}\mid\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where for any $M\in\mathbb{R}^{m\times n}$ the column and row spaces are defined as $\text{Col}(M)=\{y\in\mathbb{R}^{m}\mid\exists x\in\mathbb{R}^{n}:\;y=Mx\}$ and $\text{Row}(M)=\{y\in\mathbb{R}^{n}\mid\exists x\in\mathbb{R}^{m}:\;y=M^{\top}x\}$.
@@
Since $\Delta W=BRA$, we have
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
That is, we proved that
$\text{Col}(\Delta W)$ $\subseteq\text{Col}(B).$ (18)
Following similar arguments, one can also show $\text{Row}(\Delta W)\subseteq\text{Row}(A)$. ∎
@@ A.2 Proof of Lemma 2
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
@@
$g:=\frac{\partial L}{\partial W}\quad\&\quad g^{R}_{\text{LoRA-XS}}:=\frac{\partial L}{\partial R}.$ (19)
The chain rule gives
$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial R}\implies\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}\frac{\partial X}{\partial R}\quad\text{ for }X=RA$ (20)
@@
$\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}=sB^{\top}g\implies\frac{\partial L}{\partial R}=sB^{\top}g\frac{\partial X}{\partial R}$ (21)
Let $sB^{\top}g=y$. We know that when $X=RA$:
$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}$ (22)
[…]
@@ A.3 Proof of Theorem 3
The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (24)
@@
Since we already defined the equivalent gradient $\tilde{g}:=sBg^{R}A$, the minimization problem can be denoted as:
@@
$B^{\top}gA^{\top}=g^{R}_{\text{LoRA-XS}}/s\implies B^{\top}sBg^{R}AA^{\top}=g^{R}_{\text{LoRA-XS}}/s\implies B^{\top}Bg^{R}AA^{\top}=g^{R}_{\text{LoRA-XS}}/s^{2}$ (28)
Now since $B$ and $A$ are full rank, multiplying both sides by $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ on the left and right side respectively gives:
$(B^{\top}B)^{-1}(B^{\top}Bg^{R}AA^{\top})(AA^{\top})^{-1}$ $=(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}/s^{2}$ (29)
[…]
@@ A.4 Proof of Theorem 4
Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0$
@@
Assuming that $L$ is differentiable, we use Taylor’s theorem and get
$\Delta L$ $\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)$
$=\left\langle\frac{\partial L}{\partial R},-\eta g^{R}\right\rangle_{F}+o(\eta)$
$=-\frac{\eta}{s^{2}}\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}+o(\eta),$ (31)
where in the last step we also used the definition of $g^{R}_{\text{LoRA-XS}}$ and the result of Theorem 3. To prove $\Delta L\leq 0$ for small enough $\eta$, it is sufficient to show that
$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}\geq 0.$ (32)
Next, we note that matrices $B^{\top}B\in\mathbb{R}^{r\times r}$ and $AA^{\top}\in\mathbb{R}^{r\times r}$ are positive definite since they are positive semi-definite and matrices $B$ and $A$ are full-rank (i.e., with rank $r$) matrices, which means that $B^{\top}B$ and $AA^{\top}$ have non-zero eigenvalues. Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $U$ and $V$ such that $(B^{\top}B)^{-1}=VV^{\top}$ and $(AA^{\top})^{-1}=UU^{\top}$ (e.g., one can find such matrices using Cholesky decomposition). Then, we have
$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},VV^{\top}g^{R}_{\text{LoRA-XS}}UU^{\top}\rangle_{F}$
$=\frac{1}{s^{2}}\langle V^{\top}g^{R}_{\text{LoRA-XS}}U,V^{\top}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
$=\|V^{\top}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0.$
This concludes the proof. ∎
For our specific initialization where $(B^{\top}B)=I$, $(AA^{\top})=I$, and $s=1$, the result simplifies to:
$\Delta L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}+o(\eta)\leq 0.$ (33)
@@ A.5 Proof of Theorem 5
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A$
@@
Let $g$ be the full fine-tuning gradient. We want to prove that $\tilde{g}$ does not depend on $s$, so we try to express it in terms of $g$ which does not depend on the LoRA-XS training process or reparameterization.
@@
$g^{R}=\frac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\implies\tilde{g}=\frac{s}{s^{2}}B(B^{\top}B^{-1})g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}A$ (34)
Now since $g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}$:
$\tilde{g}=\frac{1}{s}B(B^{\top}B^{-1})sB^{\top}gA^{\top}(AA^{\top})^{-1}A=B(B^{\top}B^{-1})B^{\top}gA^{\top}(AA^{\top})^{-1}A.$ (35)
which is $s$-independent.
@@
$g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}\implies\tilde{g}=sB(sB^{\top}gA^{\top})A\implies\tilde{g}=s^{2}BB^{\top}gA^{\top}A$ (36)
which is not $s$-independent. ∎
@@ A.6 Proof of Theorem 6
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
$\Delta R=-\eta\nabla_{R}\mathcal{L}(R)\implies B\Delta RA=-\eta B\nabla_{R}\mathcal{L}(R)A.$ (37)
@@
$\Delta W$ $=-\eta\nabla_{W}\mathcal{L}(W_{0}).$ (38)
@@
$\|B\Delta RA-\eta\nabla\mathcal{L}_{W}(W_{0})\|_{F}$ $=\eta\|B\nabla_{R}\mathcal{L}(R)A-\nabla\mathcal{L}_{W}(W_{0})\|_{F}.$ (39)
@@
$\nabla_{R}\mathcal{L}$ $=B^{\top}\nabla_{W}\mathcal{L}A^{\top}.$ (40)
The problem now becomes:
$\min_{A_{\text{init}},B_{\text{init}}}\|B^{\top}(B^{\top}\nabla_{W}\mathcal{L}A^{\top})A-\nabla_{W}\mathcal{L}\|_{F}\quad\text{ where }\nabla_{W}\mathcal{L}=USV^{\top}.$ (41)
@@
$\|BB^{\top}\nabla_{W}\mathcal{L}A^{\top}A-\nabla_{W}\mathcal{L}\|_{F}$ $=\|U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}-USV^{\top}\|_{F}.$ (42)
Moreover, we also have
$U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}$ $=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}.$ (43)
The rank of $W^{\prime}$ such that
$W^{\prime}$ $=U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}$ (44)
is $\leq r$, since the corresponding ranks of $B_{\text{init}}$ and $A_{\text{init}}$ is $r$. Using the Eckart-Young Theorem, we find the optimal low-rank solution as:
$W^{\prime*}$ $=\underset{\text{rank}(W^{\prime})=r}{\text{arg min}}\|W^{\prime}-\nabla_{W}\mathcal{L}\|_{F}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}.$ (45)
Since we also get an identical expression, our solution is optimal. ∎
Appendix B Inference Overhead vs LoRA